# Mean Variance Efficient Collaborative Filtering for Stock Recommendations

### v2

@@ Abstract.
copyright: acmlicensed
@@ 1. Introduction
Recommender systems have been one of the key success factors of companies who provide online services that they could attract more customers and make them spend more time in their services (Lu et al., 2015). Numerous researchers have successfully applied recommender systems to relatively simple financial products such as bank products (Oyebode and Orji, 2020; Sharifihosseini, 2019), insurances (Qazi et al., 2017; Rokach et al., 2013; Kanchinadam et al., 2018), and real estate (Daly et al., 2014; Yuan et al., 2013). The rise of FinTech has galvanized the participation of individual investors in stock markets (Hendershott et al., 2021), but recommending stocks is not trivial due to the high uncertainty in their prices unlike other financial products.
The necessity for stock recommender systems is in no doubt since investors are known to hold under-diversified portfolios (Dimmock et al., 2020). The complexity of stock choices exacerbated by a flood of information and various market complexities often confuses individual investors. Companies offering precise and tailored stock recommendations consonant with an individual’s existing portfolio will gain distinct customer loyalty. We list three key requirements for stock recommendation as a service.
@@
In this paper, we develop a simple yet effective model for stock recommendation, called mean-variance efficient collaborative filtering (MVECF), that satisfies all three key requirements. We utilize the regularization technique to ensure that the recommendation is made based on the user’s current portfolio while increasing diversification effect by systemically handling uncertainties of stock returns, and the model is restructured to an ordinary weighted matrix factorization (WMF) form of Hu et al. (2008) to boost the computational efficiency. The experimental results show the proposed model improves the pareto optimality in a trade-off between risk and return (i.e., the mean-variance efficiency). Moreover, the model outperforms baseline stock recommendation models in recommendation accuracy and exhibits superior investment performance than the conventional recommender systems in both in-sample and out-of-sample experiments. We further show that MVECF can be easily incorporated into graph-based ranking model by applying MVECF user-item ratings to the sampling process of the ranking systems.
@@ 2.1. Collaborative Filtering
CF is the most popular approach in modern recommendation systems because of its efficiency in utilizing the entire user-item preference history. Hu et al. (2008) first developed a CF model for implicit feedback that predicts users’ preferences by minimizing the square error of preference estimations. Another type of CF would be ranking models with the Bayesian personalized ranking (BPR) objective introduced by Rendleet al. (2009). Recently, graph-based ranking models like NGCF, LightGCN, UltraGCN, and HCCF (Wang et al., 2019; He et al., 2020; Mao et al., 2021; Xia et al., 2022) have achieved strong performance and gained widespread adoption for implicit feedback data.
Our work may look similar to a stream of research on the diversity and novelty of recommendations. These involve greedy re-ranking (Ziegler et al., 2005; Zhang and Hurley, 2008; Vargas and Castells, 2014) and directly optimizing multi-objectives approaches (Su et al., 2013; Shi et al., 2012; Hurley, 2013; Wasilewski and Hurley, 2016, 2019). However, the notion of recommendation diversity and novelty is quite different from portfolio diversification. While diversity in recommender systems focus on recommending dissimilar items, and novelty seeks items that are dissimilar to the items in the user’s past experiences, portfolio diversification considers the trade-off between minimizing risk and maximizing return of the entire portfolio. For example, if we define the dissimilarity measure as negative correlation between stock returns, novelty and diversity enhancing models may recommend stocks that reduce portfolio risk, but they might easily recommend poor stocks (e.g., negative returns) just because they have low correlation with the user’s current holdings.
We incorporate LightGCN as a representative graph-based model and a model for enhancing novelty (Wasilewski and Hurley, 2019) as baseline models.
@@ 2.2. Stock Recommendations
Previous studies on stock recommender systems can be categorized into three. First, some studies focus on suggesting ways to measure similarities between stocks or equity funds and then apply existing CF methods  (Matsatsinis and Manarolis, 2009; Chang and Tu, 2018; De Rossi et al., 2020). However, simply buying stocks that are similar to the current holdings or held by similar investors may make a portfolio more exposed to a certain type of risk (e.g., sector, style), contradicting the research goal of this work. Hence, we incorporate the portfolio theory to ensure proper diversification within the CF framework.
Second, purely item-based recommendation. These studies try to find stocks that would have high returns in the future by analyzing item similarities between stocks (Yujun et al., 2016; Tu et al., 2016; Zheng et al., 2020; Gao et al., 2021). This approach has a quite different perspective from recommender systems given that most methods in this category do not utilize user information, and giving the identical recommendation to all users will never yield even minimal recommendation accuracy. In addition, accurately predicting stock returns is almost impossible, because signals are dominated by noises in financial markets. Modern Portfolio Theory (MPT), established by Harry Markowitz (1952), can recommend an items to each user that maximizes the portfolio performance when added to the user’s portfolio. We use this method as a baseline model for stock recommendation method in this category.
Third, recommendations based on user-item information and then diversification. For example, Musto et al. (2015) and Sweezy and Charron (2018) recommend stocks based on the second approach and determine weights of stocks to reduce portfolio risk. However, adjusting portfolio weights after choosing stocks would have limited effects. On the other hand, García et al. et al. (2012) and González (2012) carefully analyze user similarity and recommend a portfolio according to the user’s preference. This approach may be able to provide well-diversified portfolios to investors, but this has the same problem with existing mutual funds that investors prefer possessing a few stocks that they are interested in. In this category, the two step method of Sweezy and Charron (2018) is included in the baselines.
@@ 3.1. Weighted Matrix Factorization
Here, $y_{ui}$$\in\mathbb{R}$ is an element of Y, and $p_{u}^{T}\in{\mathbb{R}^{1\times l}}$ and $q_{i}^{T}\in{\mathbb{R}^{1\times l}}\text{\;}$ are row vectors of P and Q, respectively. $c_{ui}$$\in\mathbb{R}$ is a hyperparameter indicating the confidence level about observation $y_{ui}$}, which becomes large when ${y_{ui}}=1$ and small when ${y_{ui}}=0$, and $\lambda$ is a hyperparameter for ${\text{L}}_{2}$} regularization. $\hat{y}_{ui}$}, the estimated preference of user u to item $i$ is $p_{u}^{T}{q_{i}}$. Hu et al. (2008) proposed an efficient alternating least squares (ALS) algorithm for solving Equation (1).
@@ 3.2. Modern Portfolio Theory (MPT)
Markowitz (1952) was the first to mathematically define and analyze the risk and return of financial investments. The return of a risky asset was regarded as a random variable and the expected return was defined as its mean value and the risk was defined as its standard deviation. Then, $n$ risky assets can be described by their return vector $r\in{\mathbb{R}^{n}}$ with mean $\mu\in{\mathbb{R}^{n}}$ and covariance matrix $\Sigma\in{\mathbb{R}^{n\times n}}$. A portfolio of $n$ risky assets can be represented as a weight vector $w\in{\mathbb{R}^{n}}$, which should sum to one (i.e., $\mathop{\sum}\nolimits_{i}{w_{i}}=1\text{\;}$), and it’s expected return and risk can be expressed as ${\mu^{T}}w$ and ${w^{T}}\Sigma w$, respectively.
In Markowitz (1952), a convex quadratic programming problem (2) was proposed to find a pareto optimal portfolio between minimizing the risk and maximizing the expected return.
@@
It is called the mean-variance (MV) optimization, and the resulting optimal portfolios are called MV efficient portfolios. The term ‘efficient’ emphasizes the pareto optimality of the solution from the trade-off between risk and return, where $\gamma$ is the parameter that represents the risk aversion of the investor. Markowitz (1952) is the foundation of the MPT, and it is widely used in practice as well (Kim et al., 2021).
An evaluation metric for measuring portfolio efficiency is proposed by Sharpe (1966) which is called the Sharpe ratio. It is one of the most widely used performance measures in investment management. It is defined as a ratio between the expected return and the risk (return standard deviation) of the portfolio as
@@
It is worth noting that there are several studies that utilized the portfolio theory for information retrieval and recommender systems (Shi et al., 2012; Wasilewski and Hurley, 2017; Wang, 2009; Wang and Zhu, 2009), however, their objective is to maximize the expected relevance of item ratings (mean of ${\hat{y}}_{u}$) and minimize the risk of relevance of item ratings (variance of ${\hat{y}}_{u}$) (or some of them maximize the variance to increase recommendation diversity). Clearly, they are focusing on the mean and variance of item ratings, not the financial context (mean and variance of stock returns).
@@ 4.1. Mean-Variance Regularization
In this section, we develop the mean-variance efficient collaborative filtering (MVECF) method, which is a novel WMF model with regularization on the MV efficiency. In WMF model, the estimated item ratings (holdings) of user $u$ is ${\hat{y}_{u}}=Q{p_{u}}$, and we consider this as the user’s final portfolio. The expected return and the risk of the portfolio would be written as ${\mu^{T}}{\hat{y}_{u}}$ and $\hat{y}_{u}^{T}{\Sigma}{\hat{y}_{u}}$, respectively [^1] Notice that $y_{u}$, which is a vector of binary variables, replaces the portfolio weight $w$ in Equation 2, which should sum to one. However, we do not normalize $\hat{y}_{u}$ for simplicity. Also, most of its elements will be continue to be less than 1 during training.. We regularize ’risk – return (i.e., $\frac{\gamma}{2}\hat{y}_{u}^{T}{\Sigma}{\hat{y}_{u}}-{\mu^{T}}{\hat{y}_{u}}$)’ of the user’s portfolio, which is the objective function of the MV optimization problem in (2), to the loss function of WMF. Then, the proposed model would recommend stocks based on CF while trying to minimize risk and maximize return. The resulting formulation is given in Equation (4). $\lambda_{MV}$ is a hyperparameter that controls the trade-off between the traditional recommendation performance and the MV efficiency of recommended portfolios.
@@
If we rewrite the last MV regularization term of (4) in an elementwise expression with $\mu_{i}$} (elements of $\mu$), $\sigma_{i}^{2}$ and $\sigma_{ij}$} (diagonal and off-diagonal elements of $\Sigma$}), Equation (4) can be rewritten as Equation (5).
@@ 4.2. Restructuring MVECF into Ordinary WMF Form
Here, we further increase the computational efficiency of MVECF by restructuring it into an ordinary WMF form (as in Equation (1)) so that we can train it using the ALS algorithm developed by Hu et al. (2008). ALS is known to converge much faster than SGD, and also SGD is highly sensitive to the choice of learning rate. The trick is quite simple. In the MV regularization term in Equation (5), we change ${\hat{y}_{ui}}\mathop{\sum}\nolimits_{j:j\neq i}{\hat{y}_{uj}}{\sigma_{ij}}$ into ${\hat{y}_{ui}}\mathop{\sum}\nolimits_{j:j\neq i}{y_{uj}}{\sigma_{ij}}/\left|{{y_{u}}}\right|$. Here, $\left|{{y_{u}}}\right|$ is the number of holdings of user $u$.
@@ 5.2. Models
We consider the WMF form version of MVECF ($\text{MVECF}_{\text{wmf}}$) in Equation (9). For baseline models, we use two simple baselines (WMF and BPR) and one graph-based recommender system, LightGCN. In addition, we consider three categories of models for stock recommendations as baseline models to enhance MV efficiency of recommended portfolio with existing methods. First, the novelty enhancing BPR model ($\text{BPR}_{\text{nov}}$) of Wasilewski et al. (2019). For this model, we define the dissimilarity (distance) between two items $i$ and $j$ as $\sqrt{1-\rho_{ij}}$, where $\rho_{ij}$ is the return correlation between $i$ and $j$. Second, the 2-step method of Sweezy and Charron (2018) introduced in Section 2.1. It filters K items using a base recommendation model with a cutoff K, and then makes the final recommendation by re-ranking the scores of top-k items using the MPT method. We consider the version with WMF as the base recommendation method ($\text{2Step}_{\text{wmf}}$). Lastly, the method incorporating MPT that recommends items that maximize Sharpe ratio of recommended portfolio. We denote this method as $\text{MPT}_{\text{topSR}}$. In short, we compare $\text{MVECF}_{\text{WMF}}$ with six baseline models: WMF, BPR, LightGCN, $\text{BPR}_{\text{nov}}$, $\text{2Step}_{\text{wmf}}$, and $\text{MPT}_{\text{topSR}}$.
@@ 5.3. Performance Evaluation
As noted in Introduction, ex-post evaluation is particularly important in investment management. Hence, the same metrics, $\Delta\text{SR}$} and $\text{P}(\text{SR}>\text{SR}_{\text{init}})$, are calculated with 5 years realized return and variance of the recommended portfolio using Equation (3). Note that Sharpe ratio is undoubtedly the most widely used investment performance measure in both practice and academia.
We also analyze the improvement in mean return $\Delta\mu$} and risk $\Delta\sigma$} of recommended portfolio to see the risk-return tradeoff of MVECF. The hyperparameters are tuned within the validation set (10% of total data). Experiments are repeated with different values of balancing hyperparameter ${\lambda}_{\text{MV}}$} in 0.1, 1, 10 and risk-aversion level $\gamma$} in 1, 3, 5. More details including entire experiment results, training loss, hyperparameters of baseline models and MVECF are given in the Technical Appendix.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770427isbn: 979-8-4007-2220-2/2025/11
@@ 1. Introduction
Recommender systems have been one of the key success factors of companies who provide online services that they could attract more customers and make them spend more time in their services (Lu2015). Numerous researchers have successfully applied recommender systems to relatively simple financial products such as bank products (Oyebode2020; Sharifihosseini2019), insurances (Qazi2017; Rokach2013; Kanchinadam2018), and real estate (Daly2014; Yuan2013). The rise of FinTech has galvanized the participation of individual investors in stock markets (Hendershott2021), but recommending stocks is not trivial due to the high uncertainty in their prices unlike other financial products.
The necessity for stock recommender systems is in no doubt since investors are known to hold under-diversified portfolios (Dimmock2020). The complexity of stock choices exacerbated by a flood of information and various market complexities often confuses individual investors. Companies offering precise and tailored stock recommendations consonant with an individual’s existing portfolio will gain distinct customer loyalty. We list three key requirements for stock recommendation as a service.
@@
In this paper, we develop a simple yet effective model for stock recommendation, called mean-variance efficient collaborative filtering (MVECF), that satisfies all three key requirements. We utilize the regularization technique to ensure that the recommendation is made based on the user’s current portfolio while increasing diversification effect by systemically handling uncertainties of stock returns, and the model is restructured to an ordinary weighted matrix factorization (WMF) form of Hu et al. (Hu2008) to boost the computational efficiency. The experimental results show the proposed model improves the Pareto optimality in a trade-off between risk and return (i.e., the mean-variance efficiency). Moreover, the model outperforms baseline stock recommendation models in recommendation accuracy and exhibits superior investment performance than the conventional recommender systems in both in-sample and out-of-sample experiments. We further show that MVECF can be easily incorporated into graph-based ranking model by applying MVECF user-item ratings to the sampling process of the ranking systems.
@@ 2.1. Collaborative Filtering
CF is the most popular approach in modern recommendation systems because of its efficiency in utilizing the entire user-item preference history. Hu et al. (Hu2008) first developed a CF model for implicit feedback that predicts users’ preferences by minimizing the square error of preference estimations. Another type of CF would be ranking models with the Bayesian personalized ranking (BPR) objective introduced by Rendleet al. (Rendle2009). Recently, graph-based ranking models like NGCF, LightGCN, UltraGCN, and HCCF (Wang2019; He2020; Mao2021; Xia2022) have achieved strong performance and gained widespread adoption for implicit feedback data.
Our work may look similar to a stream of research on the diversity and novelty of recommendations. These involve greedy re-ranking (Ziegler2005; Zhang2008; Vargas2014) and directly optimizing multi-objectives approaches (Su2013; Shi2012; Hurley2013; Wasilewski2016; Wasilewski2019). However, the notion of recommendation diversity and novelty is quite different from portfolio diversification. While diversity in recommender systems focus on recommending dissimilar items, and novelty seeks items that are dissimilar to the items in the user’s past experiences, portfolio diversification considers the trade-off between minimizing risk and maximizing return of the entire portfolio. For example, if we define the dissimilarity measure as negative correlation between stock returns, novelty and diversity enhancing models may recommend stocks that reduce portfolio risk, but they might easily recommend poor stocks (e.g., negative returns) just because they have low correlation with the user’s current holdings.
We incorporate LightGCN as a representative graph-based model and a model for enhancing novelty (Wasilewski2019) as baseline models.
@@ 2.2. Stock Recommendations
Previous studies on stock recommender systems can be categorized into three. First, some studies focus on suggesting ways to measure similarities between stocks or equity funds and then apply existing CF methods  (Matsatsinis2009; Chang2018; DeRossi2020). However, simply buying stocks that are similar to the current holdings or held by similar investors may make a portfolio more exposed to a certain type of risk (e.g., sector, style), contradicting the research goal of this work. Hence, we incorporate the portfolio theory to ensure proper diversification within the CF framework.
Second, purely item-based recommendation. These studies try to find stocks that would have high returns in the future by analyzing item similarities between stocks (Yujun2016; Tu2016; Zheng2020; Gao2021). This approach has a quite different perspective from recommender systems given that most methods in this category do not utilize user information, and giving the identical recommendation to all users will never yield even minimal recommendation accuracy. In addition, accurately predicting stock returns is almost impossible, because signals are dominated by noises in financial markets. Modern Portfolio Theory (MPT), established by Harry Markowitz (Markowitz1952), can recommend an items to each user that maximizes the portfolio performance when added to the user’s portfolio. We use this method as a baseline model for stock recommendation method in this category.
Third, recommendations based on user-item information and then diversification. For example, Musto et al. (Musto2015) and Sweezy and Charron (Sweezy2018) recommend stocks based on the second approach and determine weights of stocks to reduce portfolio risk. However, adjusting portfolio weights after choosing stocks would have limited effects. On the other hand, García et al. (Garcia2012) and González (Gonzalez2012) carefully analyze user similarity and recommend a portfolio according to the user’s preference. This approach may be able to provide well-diversified portfolios to investors, but this has the same problem with existing mutual funds that investors prefer possessing a few stocks that they are interested in. In this category, the two step method of Sweezy and Charron (Sweezy2018) is included in the baselines.
@@ 3.1. Weighted Matrix Factorization
Here, $y_{ui}$$\in\mathbb{R}$ is an element of Y, and $p_{u}^{T}\in{\mathbb{R}^{1\times l}}$ and $q_{i}^{T}\in{\mathbb{R}^{1\times l}}\text{\;}$ are row vectors of P and Q, respectively. $c_{ui}$$\in\mathbb{R}$ is a hyperparameter indicating the confidence level about observation $y_{ui}$}, which becomes large when ${y_{ui}}=1$ and small when ${y_{ui}}=0$, and $\lambda$ is a hyperparameter for ${\text{L}}_{2}$} regularization. $\hat{y}_{ui}$}, the estimated preference of user u to item $i$ is $p_{u}^{T}{q_{i}}$. Hu et al. (Hu2008) proposed an efficient alternating least squares (ALS) algorithm for solving Equation (1).
@@ 3.2. Modern Portfolio Theory (MPT)
Markowitz (Markowitz1952) was the first to mathematically define and analyze the risk and return of financial investments. The return of a risky asset was regarded as a random variable and the expected return was defined as its mean value and the risk was defined as its standard deviation. Then, $n$ risky assets can be described by their return vector $r\in{\mathbb{R}^{n}}$ with mean $\mu\in{\mathbb{R}^{n}}$ and covariance matrix $\Sigma\in{\mathbb{R}^{n\times n}}$. A portfolio of $n$ risky assets can be represented as a weight vector $w\in{\mathbb{R}^{n}}$, which should sum to one (i.e., $\mathop{\sum}\nolimits_{i}{w_{i}}=1\text{\;}$), and it’s expected return and risk can be expressed as ${\mu^{T}}w$ and $w^{T}\Sigma w$, respectively.
In Markowitz (Markowitz1952), a convex quadratic programming problem (2) was proposed to find a Pareto optimal portfolio between minimizing the risk and maximizing the expected return.
@@
It is called the mean-variance (MV) optimization, and the resulting optimal portfolios are called MV efficient portfolios. The term ‘efficient’ emphasizes the Pareto optimality of the solution from the trade-off between risk and return, where $\gamma$ is the parameter that represents the risk aversion of the investor. Markowitz (Markowitz1952) is the foundation of the MPT, and it is widely used in practice as well (Kim2021).
An evaluation metric for measuring portfolio efficiency is proposed by Sharpe (Sharpe1966) which is called the Sharpe ratio. It is one of the most widely used performance measures in investment management. It is defined as a ratio between the expected return and the the standard deviation of portfolio returns as
@@
It is worth noting that there are several studies that utilized the portfolio theory for information retrieval and recommender systems (Shi2012; Wasilewski2017; Wang2009; Wang2009_2), however, their objective is to maximize the expected relevance of item ratings (mean of ${\hat{y}}_{u}$) and minimize the risk of relevance of item ratings (variance of ${\hat{y}}_{u}$) (or some of them maximize the variance to increase recommendation diversity). Clearly, they are focusing on the mean and variance of item ratings, not the financial context (mean and variance of stock returns).
@@ 4.1. Mean-Variance Regularization
In this section, we develop the mean-variance efficient collaborative filtering (MVECF) method, which is a novel WMF model with regularization on the MV efficiency. In WMF model, the estimated item ratings (holdings) of user $u$ is ${\hat{y}_{u}}=Q{p_{u}}$, and we consider this as the user’s final portfolio. The expected return and the risk of the portfolio would be written as ${\mu^{T}}{\hat{y}_{u}}$ and $\hat{y}_{u}^{T}{\Sigma}{\hat{y}_{u}}$, respectively [^1] Notice that $y_{u}$, which is a vector of binary variables, replaces the portfolio weight $w$ in Equation 2, which should sum to one. However, we do not normalize $\hat{y}_{u}$ for simplicity. Also, most of its elements will be continue to be less than 1 during training.. We regularize ‘risk – return (i.e., $\frac{\gamma}{2}\hat{y}_{u}^{T}{\Sigma}{\hat{y}_{u}}-{\mu^{T}}{\hat{y}_{u}}$)’ of the user’s portfolio, which is the objective function of the MV optimization problem in (2), to the loss function of WMF. Then, the proposed model would recommend stocks based on CF while trying to minimize risk and maximize return. The resulting formulation is given in Equation (4). $\lambda_{MV}$ is a hyperparameter that controls the trade-off between the traditional recommendation performance and the MV efficiency of recommended portfolios.
@@
If we rewrite the last MV regularization term of Equation (4) in an elementwise expression with $\mu_{i}$} (elements of $\mu$), $\sigma_{i}^{2}$ and $\sigma_{ij}$} (diagonal and off-diagonal elements of $\Sigma$}), Equation (4) can be rewritten as Equation (5).
@@ 4.2. Restructuring MVECF into Ordinary WMF Form
Here, we further increase the computational efficiency of MVECF by restructuring it into an ordinary WMF form (as in Equation (1)) so that we can train it using the ALS algorithm developed by Hu et al. (Hu2008). ALS is known to converge much faster than SGD, and also SGD is highly sensitive to the choice of learning rate. The trick is quite simple. In the MV regularization term in Equation (5), we change ${\hat{y}_{ui}}\mathop{\sum}\nolimits_{j:j\neq i}{\hat{y}_{uj}}{\sigma_{ij}}$ into ${\hat{y}_{ui}}\mathop{\sum}\nolimits_{j:j\neq i}{y_{uj}}{\sigma_{ij}}/\left|{{y_{u}}}\right|$. Here, $\left|{{y_{u}}}\right|$ is the number of holdings of user $u$.
@@ 5.2. Models
We consider the WMF form version of MVECF ($\text{MVECF}_{\text{wmf}}$) in Equation (9). For baseline models, we use two simple baselines (WMF and BPR) and one graph-based recommender system, LightGCN. In addition, we consider three categories of models for stock recommendations as baseline models to enhance MV efficiency of recommended portfolio with existing methods. First, the novelty enhancing BPR model ($\text{BPR}_{\text{nov}}$) of Wasilewski et al. (Wasilewski2019). For this model, we define the dissimilarity (distance) between two items $i$ and $j$ as $\sqrt{1-\rho_{ij}}$, where $\rho_{ij}$ is the return correlation between $i$ and $j$. Second, the 2-step method of Sweezy and Charron (Sweezy2018) introduced in Section 2.1. It filters K items using a base recommendation model with a cutoff K, and then makes the final recommendation by re-ranking the scores of top-k items using the MPT method. We consider the version with WMF as the base recommendation method ($\text{2Step}_{\text{wmf}}$). Lastly, the method incorporating MPT that recommends items that maximize Sharpe ratio of recommended portfolio. We denote this method as $\text{MPT}_{\text{topSR}}$. In short, we compare $\text{MVECF}_{\text{WMF}}$ with six baseline models: WMF, BPR, LightGCN, $\text{BPR}_{\text{nov}}$, $\text{2Step}_{\text{wmf}}$, and $\text{MPT}_{\text{topSR}}$.
@@ 5.3. Performance Evaluation
As noted in Introduction, ex-post evaluation is particularly important in investment management. Hence, the same metrics, $\Delta\text{SR}$} and $\text{P}(\text{SR}>\text{SR}_{\text{init}})$, are calculated with 5 years realized return and risk of the recommended portfolio using Equation (3). Note that Sharpe ratio is undoubtedly the most widely used investment performance measure in both practice and academia.
We also analyze the improvement in mean return $\Delta\mu$} and risk $\Delta\sigma$} of recommended portfolio to see the risk-return tradeoff of MVECF. The hyperparameters are tuned within the validation set (10% of total data). Experiments are repeated with different values of balancing hyperparameter ${\lambda}_{\text{MV}}$} in 0.1, 1, 10 and risk-aversion level $\gamma$} in 1, 3, 5.